# What is the Curvature of 2D Euclidean Quantum Gravity?

### v1

@@ 1 Introduction
The remainder of the paper is organized as follows. In Sec. 2, we discuss quantum-gravitational observables and recall the notion of the quantum Ricci curvature and why it is physically important. Sec. 3 introduces the three different ensembles of 2D triangulations in terms of the local geometric degeneracies they allow for. Sec. 4 contains a description of the theoretical and numerical set-up of our investigation. In Sec. 5 we present our main numerical results, starting with a series of Hausdorff dimension measurements to test the simulation set-up and get a first idea of the influence of the choice of ensemble. This is followed by a detailed investigation of the curvature profile, where we critically analyze and extend the methodology and results of [9], and the role of lattice effects in particular. We conclude in Sec. 6 with a summary and outlook.
@@ 2 Observables and quantum Ricci curvature
Fig. 2 shows $\bar{d}_{\mathrm{av}}(\delta)/\delta\!\equiv\!\bar{d}(\delta)/\delta$ for continuum spheres of dimension $D\!=\!$ 2, 3, 4 and 5. Without loss of generality, the distance $\delta$ is given in units of $\rho$, the curvature radius of the spheres. In line with our earlier remarks on the nature of the quantum Ricci curvature, all curves decrease initially, reflecting the presence of positive Ricci curvature. They start at different constant values, which increase with the dimension, but are otherwise qualitatively similar. Beyond $\delta\!\approx\!\rho$, the curves coincide within the plot’s resolution. Before discussing the measurements of the curvature profile in the quantum theory, we will describe next the different geometric ensembles for which these numerical experiments will be performed.
@@ 3 The three geometric ensembles
In full 4D quantum gravity these issues are the subject of ongoing research [24, 2, 6], but fortunately the situation in 2D is much simpler, because of the absence of local physical degrees of freedom and UV-divergences, and the analytical control we have over many aspects of the theory. There are two known universality classes for pure gravity models without matter [^5] We also do not consider models which include summing over topologies or allow for topology changes as a function of time., characterized by a set of critical exponents. Relevant for this paper is the universality class of 2D Euclidean quantum gravity, which is characterized by a Hausdorff dimension $d_{H}\!=\!4$, spectral dimension $d_{S}\!=\!2$ and entropy exponent $\gamma\!=\!-1/2$, and contains Liouville quantum gravity for $\gamma\!=\!\sqrt{8/3}$ [25, 26], 2D Euclidean DT [27, 28] and the Brownian sphere [29, 30]. The other universality class is that of Lorentzian or causal models, characterized by $d_{H}\!=\!2$, $d_{S}\!=\!2$ and $\gamma\!=\!1/2$ and contains 2D gravity in proper-time gauge [31], 2D CDT [32] and the uniform infinite planar tree [33].
@@
The standard choice of configuration space for gravitational Euclidean path integrals in a DT formulation are so-called simplicial manifolds [34], also known as combinatorial triangulations [35] or combinatorial manifolds [36], whose definition we will give below. The associated combinatorial ensemble of such objects will be denoted by ${\cal T}_{c}$. Relevant for our investigation is the existence in 2D of two natural generalized ensembles, the restricted degenerate ensemble ${\cal T}_{r}$ and the maximally degenerate ensemble ${\cal T}_{m}$, with ${\cal T}_{c}\!\subset\!{\cal T}_{r}\!\subset\!{\cal T}_{m}$, whose geometric configurations are allowed to be locally less regular in ways we will specify below. [^6] We follow the nomenclature of [35] for the different ensembles. They beautifully illustrate the concept of universality, since they are known to lead to the same continuum limit of Euclidean 2D quantum gravity [37]. Here we will exploit the fact that on finite lattices the different ensembles behave differently with regard to finite-size and lattice discretization effects (see e.g. [35] and references therein). However, what represents a best choice depends on the observable under study and can be difficult to predict beforehand. It turns out that for the curvature profile introduced in Sec. 2, we obtain new insights by numerically simulating it on all three ensembles. As we will demonstrate in Sec. 5.2, this allows us to draw sharper conclusions on the comparison with continuum spheres and improve on earlier results [9].
There are several ways to characterize the local geometry of configurations in the three ensembles. Our starting point are finite triangulations with the topology of a two-sphere [^7] The notation $S^{(d)}$ will be used for $d$-spheres as topological spaces, to distinguish them from the spheres $S^{\delta}$ of radius $\delta$ defined in eq. (4) as subsets of vertices on a given triangulation. $S^{(2)}$, whose elementary building blocks are flat equilateral triangles (two-simplices $\sigma^{(2)}$) of edge length $a$, the already mentioned UV-cutoff. $D$-simplices $\sigma^{(D)}$ of lower dimension $D$ contained in the triangles are their edges or links (one-simplices $\sigma^{(1)}$) and vertices (zero-simplices $\sigma^{(0)}$). The ensembles differ by how the triangles can be glued together pairwise along their edges to obtain a piecewise flat geometry $T$.
In what follows we will need a few definitions regarding simplicial structures [34, 36]. A face of a simplex $\sigma$ is any simplex whose vertices are a subset of the vertices of $\sigma$. The star st$(\sigma)$ of a simplex $\sigma$ in a triangulation $T$ is the union of all simplices in $T$ that have $\sigma$ as a face. Lastly, the link lk$(\sigma)$ of a simplex $\sigma$ is the union of all simplices $\tau\!\in\!\mathrm{st}(\sigma)$ which have vanishing intersection with $\sigma$, $\tau\cap\sigma\!=\!\emptyset$. Fig. 3 shows an example of the star and link of a zero-simplex $\sigma^{(0)}$. [^8] The notion of “link” introduced here should not be confused with what elsewhere we use as a synonym for “edge”.
A triangulation $T$ of $S^{(2)}$ is combinatorial, $T\!\in\!{\cal T}_{c}$, if each of its triangles and edges is combinatorially distinct. By this we mean that each triangle is defined uniquely by a set of three distinct vertices, and each edge by a set of two distinct vertices. This excludes e.g. the presence of two distinct edges that share the same vertices, or of two distinct triangles that have more than one edge in common. The graph dual to a 2D combinatorial triangulation is a trivalent $\phi^{3}$-graph without tadpole or self-energy subgraphs [35]. Yet another way to characterize elements of ${\cal T}_{c}$ uses the definitions just introduced: a triangulation is combinatorial if the link of each zero-simplex is homeomorphic to a one-sphere (cf. Fig. 3), and the link of each one-simplex is homeomorphic to a zero-sphere (the set of two points),
@@ 4 Theoretical and numerical set-up
where $\Lambda$ denotes the cosmological constant. The functional integration is over the space of geometries $[g]$ (metrics modulo diffeomorphisms) on a given manifold $M$, in our case a two-sphere $S^{(2)}$. The action $S$ is the Einstein-Hilbert action, which in two dimensions reduces to a cosmological-constant term, since the integrated Ricci scalar is a topological term which drops out when computing expectation values.
@@
where the sums are over triangulations of a given discrete volume $N_{2}$. The fixed-volume path integral $Z_{N_{2}}$ in (13) is related to the DT path integral (11) for cosmological constant $\lambda$ by a Laplace transform,
@@ 5.1 Hausdorff dimension
for small $r$. We noted earlier that the point set $S^{r}_{p}$ at a vertex $p\!\in\!T$ defined by eq. (4) in general does not form a true sphere, in the sense that the edges between pairs of neighbouring vertices do not line up to form a $S^{(1)}$ topologically. For $r\!=\!1$, we have already seen examples of this for $S^{1}_{p}$, with $p\!=\!{\sigma^{(0)}_{2}}$, in Figs. 5 and 7. Similar phenomena are present for larger $r$, also in the combinatorial ensemble. The overbar in eq. (16) denotes a configuration average, which for a local geometric quantity ${\cal O}(p)$ on a triangulation $T$ is defined by
@@
$\rho_{N_{2}}(r)=P_{4}(r)\exp(-ar^{b}),$ (22)
for a fourth-order polynomial $P_{4}(r)$ with associated fitting coefficients and two additional fitting parameters $a$ and $b$. The maximum is then found by using the fitted function, and the locations and heights of the maxima for different $N_{2}$ can subsequently be fitted to (21). The estimates for $d_{H}$ we obtained from this procedure are listed in the first two data columns of Table 1. We observe that for both methods the Hausdorff dimensions measured in the combinatorial and restricted degenerate ensembles significantly underestimate the analytically known value of $d_{H}\!=\!4$. This is a familiar feature from the long history of measurements of this observable in 2D EDT (see e.g. the discussion in [44]). Our measurements show a clear improvement when using a less regular lattice, presumably because lattices that allow for a smaller coordination number (number $n$ of edges meeting at a vertex) explore the local geometry more effectively and reduce finite-size effects. [^10] The minimal coordination number is $n_{\mathrm{min}}\!=\!3$, 2 and 1 in the ensembles ${\cal T}_{c}$, ${\cal T}_{r}$ and ${\cal T}_{m}$ respectively. The authors of [44] worked with the maximally degenerate ensemble ${\cal T}_{m}$, took data for six different lattice sizes in the range $N_{2}\!\in\![1\mathrm{k},32\mathrm{k}]$, and reported $d_{H}\!=\!$3.835(59)$$ from peak location and $d_{H}\!=\!$4.040(98)$$ from peak height measurements. Comparing these values with our results, they lie just outside each others’ statistical error bars, which we take as an indication that they are broadly compatible.
@@
Lastly, we used a variant of the previous method, which was introduced in [45], where again entire curves are brought to overlap. The difference is that one looks for an optimal collapse to a specific reference curve, given by the distribution $\rho_{N_{2}}(x)$ for the maximal volume, in our case $N_{2,\mathrm{max}}\!=\!1000$k. Also this method uses a constant shift $s$ of the argument $x$ for a given ensemble $\cal T$ to obtain a best curve collapse. The quality of the collapses we found with this method [^11] Due to insufficient statistics, we had to omit a subleading term in $N_{2}/N_{2,\mathrm{max}}$ used in [45]. were very good, especially for the degenerate ensembles (see [23] for further details), and resulted in the Hausdorff dimension estimates collected in the last column of Table 1. They follow the pattern of the other columns in that the maximally degenerate ensemble gives the dimension closest to the correct value of 4.121212We do not have a clear explanation for why the dimensions in the last column do not grow monotonically; a closer examination of this issue lies beyond the scope of this paper. The sizes of the statistical error bars are smaller than those for the other methods, but the actual estimates of $d_{H}$ still do not lie within a standard deviation of $d_{H}\!=\!4$, but are systematically lower, presumably due to finite-size effects. The work of [45], which used quadrangulations instead of triangulations, has shown that a dedicated numerical effort and much larger lattices (of up to $172^{24}\!\approx\!$ 17m squares) are needed to do significantly better, in that case yielding a “precision estimate” of $d_{H}\!=\!$3.9970(13)$$.
@@ 5.2.1 Analysis of the combinatorial ensemble reloaded
One feature that remains to be addressed before turning to the measurements is how to identify the lattice analogue of the $\delta$-independent constant $c_{\mathrm{av}}$ on the right-hand side of eq. (8). This is important when comparing to the curvature profile of continuum spaces, in our case round spheres of dimension $D$ and curvature radius $\rho$, because the value of this constant is nonuniversal and depends on the discretization [1]. Due to short-distance lattice artefacts, defining $c_{\mathrm{av}}$ by a limit $\delta\!\rightarrow\!0$, as in the smooth classical case, is not a meaningful prescription. Instead, one usually sets $c_{\mathrm{av}}\!:=\!\langle\bar{d}_{\mathrm{av}}(\delta)/\delta\rangle|_{\delta=\delta_{\mathrm{min}}}$, where $\delta_{\mathrm{min}}$ is chosen such that for $\delta\!\geq\!\delta_{\mathrm{min}}$ the measured data are considered to be free of significant discretization artefacts.
@@
Although the (average) triangulations are clearly not exact spheres globally, as can be seen from the asymmetry of the curves for the shell volumes in Fig. 9, the scaling ansatz (24) does not depend on these details and should be rather robust. Following [9], this motivates the assumption that for a given choice of $D\!=\!2$, 3, 4 or 5, the discrete volume $N_{2}$ scales with the $\cal D$th power of the effective curvature radius,
@@
The outcomes are collected in Table 2, next to the corresponding results found in reference [9]. The latter are systematically smaller than ours by at least one standard deviation. Since details of the derivation of [9] were no longer available, we performed multiple cross-checks of our implementation in search of possible errors, but did not find any. The normalization error mentioned in footnote 13 would lead to smaller values of $\cal D$, like those reported in [9], and may explain at least part of the discrepancy with our estimates [23]. Despite leaving this issue open, our conclusion from the results in Table 2 for the combinatorial ensemble is identical to that reached in [9], namely that $\cal D$ and $D$ match best when $D\!=\!5$.
However, repeating the same set of measurements on the degenerate ensembles ${\cal T}_{r}$ and ${\cal T}_{m}$ throws new light on the (non-)reliability of the data for small $N_{2}$ measured in the ${\cal T}_{c}$-ensemble. For illustration, we again show log-log plots of the effective curvature radius $\rho_{\mathrm{eff}}$ as a function of the volume $N_{2}$ for the case $D\!=\!4$ (Fig. 12), while the best sphere fitting dimensions ${\cal D}$ for all $D$ are collected in Table 3. Similar to the results for the combinatorial ensemble, $\cal D$ decreases with increasing $D$, and both dimensions match best at $D\!=\!5$. Before taking this as a confirmation of the conclusion of [9], we note that there is a significant difference with the combinatorial ensemble. In contrast to the apparently good fit of Fig. 11, it is clear from Fig. 12 that for the other two ensembles a power law is not a good fit for the data, although our Hausdorff dimension measurements suggested that these ensembles should suffer less from lattice effects. We have observed this behaviour not only for $D\!=\!4$, but also for the other values of $D$. Roughly speaking, the slope is flatter for the three smallest triangulation sizes, which may signal the presence of finite-size effects. Alerted by these findings, we have performed a detailed analysis of discretization and finite-size effects, which will be presented in the following subsections.

### v2

@@ 1 Introduction
The remainder of the paper is organized as follows. In Sec. 2, we discuss quantum-gravitational observables, recalling the notion of quantum Ricci curvature and why it is physically important. Sec. 3 introduces the three different ensembles of 2D triangulations. Sec. 4 contains a description of the theoretical and numerical set-up of our investigation. In Sec. 5 we present our main numerical results, starting with a series of Hausdorff dimension measurements to test the simulation set-up and get a first idea of the influence of the choice of ensemble. This is followed by an investigation of the curvature profile, where we critically analyze and extend the methodology and results of [9], and the role of lattice effects in particular. We conclude in Sec. 6 with a summary and outlook. The Appendix contains details of the local geometry of triangulations in the different ensembles.
@@ 2 Observables and quantum Ricci curvature
Fig. 2 shows $\bar{d}_{\mathrm{av}}(\delta)/\delta\!\equiv\!\bar{d}(\delta)/\delta$ for continuum spheres of dimension $D\!=\!$ 2, 3, 4 and 5. Without loss of generality, the distance $\delta$ is given in units of $\rho$, the curvature radius of the spheres. In line with our earlier remarks on the nature of the quantum Ricci curvature, all curves decrease initially, reflecting the presence of positive Ricci curvature. They start at different constant values, which increase with the dimension, but are otherwise qualitatively similar. Beyond $\delta\!\approx\!\rho$, the curves coincide within the plot’s resolution. Before discussing the measurements of the curvature profile in the quantum theory, we introduce next the different geometric ensembles for which these numerical experiments will be performed.
@@ 3 The three geometric ensembles
In full 4D quantum gravity these issues are the subject of ongoing research [24, 2, 6], but fortunately the situation in 2D is much simpler, because of the absence of local physical degrees of freedom and UV-divergences, and the analytical control we have over many aspects of the theory. There are two known universality classes for pure gravity models without matter [^5] We also do not consider models which include summing over topologies or allow for topology changes as a function of time., characterized by a set of critical exponents. Relevant for this paper is the universality class of 2D Euclidean quantum gravity, which is characterized by a Hausdorff dimension $d_{H}\!=\!4$, spectral dimension $d_{S}\!=\!2$ and entropy exponent $\gamma\!=\!-1/2$, and contains Liouville quantum gravity for $\gamma^{L}\!=\!\sqrt{8/3}$ [25, 26], 2D Euclidean DT [27, 28] and the Brownian sphere [29, 30]. The other universality class is that of Lorentzian or causal models, characterized by $d_{H}\!=\!2$, $d_{S}\!=\!2$ and $\gamma\!=\!1/2$ and contains 2D gravity in proper-time gauge [31], 2D CDT [32] and the uniform infinite planar tree [33].
The standard choice of configuration space for gravitational Euclidean path integrals in a DT formulation are so-called simplicial manifolds [34], also known as combinatorial triangulations [35] or combinatorial manifolds [36], whose definition will be given below. The associated combinatorial ensemble of such objects will be denoted by ${\cal T}_{c}$. Relevant for our investigation is the existence in 2D of two natural generalized ensembles, the restricted degenerate ensemble ${\cal T}_{r}$ and the maximally degenerate ensemble ${\cal T}_{m}$, with ${\cal T}_{c}\!\subset\!{\cal T}_{r}\!\subset\!{\cal T}_{m}$, whose geometric configurations are allowed to be locally less regular. [^6] We follow the nomenclature of [35] for the different ensembles. The most straightforward way to characterize them is in terms of their dual, trivalent planar graphs: ${\cal T}_{m}$ corresponds to general such graphs, ${\cal T}_{r}$ to graphs without tadpoles, and ${\cal T}_{c}$ to graphs without tadpole or self-energy subgraphs. Equivalent characterizations of these ensembles and their irregularities in terms of the local geometry of the triangulations and the combinatorial data used in the simulations are described in the Appendix.
@@
The three ensembles beautifully illustrate the concept of universality, since they are known to lead to the same continuum limit of Euclidean 2D quantum gravity. [^7] How this universality is captured in a matrix model formulation is discussed in [37]. Here we will exploit the fact that on finite lattices the different ensembles behave differently with regard to finite-size and lattice discretization effects (see e.g. [35] and references therein). However, what represents a best choice depends on the observable under study and can be difficult to predict beforehand. It turns out that for the curvature profile introduced in Sec. 2, we obtain new insights by numerically simulating it on all three ensembles. As we will demonstrate in Sec. 5.2, this allows us to draw sharper conclusions on the comparison with continuum spheres and improve on earlier results [9].
@@ 4 Theoretical and numerical set-up
where $\Lambda$ denotes the cosmological constant. The functional integration is over the space of geometries $[g]$ (metrics modulo diffeomorphisms) on a given manifold $M$, in our case a two-sphere [^8] The notation $S^{(d)}$ will be used for $d$-spheres as topological spaces, to distinguish them from the spheres $S^{\delta}$ of radius $\delta$ defined in eq. (4) as subsets of vertices on a given triangulation. $S^{(2)}$. The action $S$ is the Einstein-Hilbert action, which in two dimensions reduces to a cosmological-constant term, since the integrated Ricci scalar is a topological term which drops out when computing expectation values.
@@
where the sums are over triangulations of a given discrete volume $N_{2}$. The fixed-volume partition function $Z_{N_{2}}$ in (12) is related to the DT path integral (10) for cosmological constant $\lambda$ by a Laplace transform,
@@ 5.1 Hausdorff dimension
for small $r$. We noted earlier that the point set $S^{r}_{p}$ at a vertex $p\!\in\!T$ defined by eq. (4) in general does not form a true sphere, in the sense that the edges between pairs of neighbouring vertices do not line up to form a $S^{(1)}$ topologically. For $r\!=\!1$, examples of this for $S^{1}_{p}$, with $p\!=\!{\sigma^{(0)}_{2}}$, are given in Figs. 16 and 18 of the appendix. Similar phenomena are present for larger $r$, also in the combinatorial ensemble. The overbar in eq. (15) denotes a configuration average, which for a local [^10] Here and elsewhere, “(quasi-)local” should be understood in a lattice sense; due to the discretization, a (quasi-)local quantity associated with a vertex $p$ will typically depend on fields not just at $p$ but also at (sub-)simplices in a small neighbourhood of $p$. geometric quantity ${\cal O}(p)$ on a triangulation $T$ is defined by
@@
$\rho_{N_{2}}(r)=P_{4}(r)\exp(-\alpha r^{\beta}),$ (21)
for a fourth-order [^11] According to [44], $l\!=\!4$ is the lowest order of a polynomial $P_{l}$ that leads to a “reasonably good” fit with the ansatz (21). polynomial $P_{4}(r)$ with associated fitting coefficients and two additional fitting parameters $\alpha$ and $\beta$. The maximum is then found by using the fitted function, and the locations and heights of the maxima for different $N_{2}$ can subsequently be fitted to (20). The estimates for $d_{H}$ we obtained from this procedure are listed in the first two data columns of Table 1. We observe that for both methods the Hausdorff dimensions measured in the combinatorial and restricted degenerate ensembles significantly underestimate the analytically known value of $d_{H}\!=\!4$. This is a familiar feature from the long history of measurements of this observable in 2D EDT (see e.g. the discussion in [44]). Our measurements show a clear improvement when using a less regular lattice, presumably because lattices that allow for a smaller coordination number (number $n$ of edges meeting at a vertex) explore the local geometry more effectively and reduce finite-size effects. [^12] The minimal coordination number is $n_{\mathrm{min}}\!=\!3$, 2 and 1 in the ensembles ${\cal T}_{c}$, ${\cal T}_{r}$ and ${\cal T}_{m}$ respectively. The authors of [44] worked with the maximally degenerate ensemble ${\cal T}_{m}$, took data for six different lattice sizes in the range $N_{2}\!\in\![1\mathrm{k},32\mathrm{k}]$, and reported $d_{H}\!=\!$3.835(59)$$ from peak location and $d_{H}\!=\!$4.040(98)$$ from peak height measurements. Comparing these values with our results, they lie just outside each others’ statistical error bars, which we take as an indication that they are broadly compatible.
@@
Lastly, we used a variant of the previous method, which was introduced in [45], where again entire curves are brought to overlap. The difference is that one looks for an optimal collapse to a specific reference curve, given by the distribution $\rho_{N_{2}}(x)$ for the maximal volume, in our case $N_{2,\mathrm{max}}\!=\!1000$k. Also this method uses a constant shift $s$ of the argument $x$ for a given ensemble $\cal T$ to obtain a best curve collapse. The quality of the collapses we found with this method [^13] Due to insufficient statistics, we had to omit a subleading term in $N_{2}/N_{2,\mathrm{max}}$ used in [45]. were very good, especially for the degenerate ensembles (see [23] for further details), and resulted in the Hausdorff dimension estimates collected in the last column of Table 1. They follow the pattern of the other columns in that the maximally degenerate ensemble gives the dimension closest to the correct value of 4.141414We do not have a clear explanation for why the dimensions in the last column do not grow monotonically; a closer examination of this issue lies beyond the scope of this paper. The sizes of the statistical error bars are smaller than those for the other methods, but the actual estimates of $d_{H}$ still do not lie within a standard deviation of $d_{H}\!=\!4$, but are systematically lower, presumably due to finite-size effects. The work of [45], which used quadrangulations instead of triangulations, has shown that a dedicated numerical effort and much larger lattices (of up to $172^{24}\!\approx\!$ 17m squares) are needed to do significantly better, in that case yielding a “precision estimate” of $d_{H}\!=\!$3.9970(13)$$.
@@ 5.2.1 Analysis of the combinatorial ensemble reloaded
One feature that remains to be addressed before turning to the measurements is how to identify the lattice analogue of the $\delta$-independent constant $c_{\mathrm{av}}$ on the right-hand side of eq. (8). This is important when comparing to the curvature profile of continuum spaces, in our case round spheres of dimension $D$ and curvature radius $\rho$, because the value of this constant is nonuniversal and depends on the discretization [1]. Due to short-distance lattice artefacts, defining $c_{\mathrm{av}}$ by a limit $\delta\!\rightarrow\!0$, as was done in the smooth classical case, is not a meaningful prescription. It should instead be chosen at a value $\delta_{\mathrm{min}}$ as close as possible to $\delta\!=\!0$, but outside the small-$\delta$ region dominated by discretization artefacts, defining $c_{\mathrm{av}}\!:=\!\langle\bar{d}_{\mathrm{av}}(\delta)/\delta\rangle|_{\delta=\delta_{\mathrm{min}}}$. The optimal choice of $\delta_{\mathrm{min}}$ and how different choices affect the final results should be examined on a case-by-case basis. For two-dimensional triangulations studied so far, including regular lattices and Delaunay triangulations [1], and configurations of DT [9] and CDT quantum gravity [46, 21], significant lattice artefacts have been found in the region $\delta\lesssim\delta_{\mathrm{min}}\!\approx\!5$. As we will see in Sec. 5.2.2 below, the different ensembles used in our re-examination of the curvature profiles also require a fresh look at $\delta_{\mathrm{min}}$.
@@
Although the (average) triangulations are clearly not exact spheres globally, as can be seen from the asymmetry of the curves for the shell volumes in Fig. 4, the scaling ansatz (23) does not depend on these details and should be rather robust. Following [9], we assume that the quantum geometry resembles a continuum $\cal D$-sphere globally, in the sense that its volume scales with the ${\cal D}$th power of its curvature radius. The consistency of this (mild) assumption is shown below. It implies that for a given choice of $D\!=\!2$, 3, 4 or 5, the discrete volume $N_{2}$ is expected to scale with the $\cal D$th power of the effective curvature radius,
@@
The outcomes are collected in Table 2, next to the corresponding results found in reference [9]. The latter are systematically smaller than ours by at least one standard deviation. Since details of the derivation of [9] were no longer available, we performed multiple cross-checks of our implementation in search of possible errors, but did not find any. The normalization error mentioned in footnote 15 would lead to smaller values of $\cal D$, like those reported in [9], and may explain at least part of the discrepancy with our estimates [23]. Despite leaving this issue open, our conclusion from the results in Table 2 for the combinatorial ensemble is identical to that reached in [9], namely that $\cal D$ and $D$ match best when $D\!=\!5$.
@@
However, repeating the same set of measurements on the degenerate ensembles ${\cal T}_{r}$ and ${\cal T}_{m}$ throws new light on the (non-)reliability of the data for small $N_{2}$ measured in the ${\cal T}_{c}$-ensemble. For illustration, we again show log-log plots of the effective curvature radius $\rho_{\mathrm{eff}}$ as a function of the volume $N_{2}$ for the case $D\!=\!4$ (Fig. 7), while the best sphere fitting dimensions ${\cal D}$ for all $D$ are collected in Table 3.
@@
Similar to the results for the combinatorial ensemble, $\cal D$ decreases with increasing $D$, and both dimensions match best at $D\!=\!5$. Before taking this as a confirmation of the conclusion of [9], we note that there is a significant difference with the combinatorial ensemble. In contrast to the apparently good fit of Fig. 6, it is clear from Fig. 7 that for the other two ensembles a power law is not a good fit for the data, although our Hausdorff dimension measurements suggested that these ensembles should suffer less from lattice effects. We have observed this behaviour not only for $D\!=\!4$, but also for the other values of $D$. Roughly speaking, the slope is flatter for the three smallest triangulation sizes, which may signal the presence of finite-size effects. Alerted by these findings, we have performed a detailed analysis of discretization and finite-size effects, which will be presented in the following subsections.